OpenStudy (joshoyen):

Find the absolute value of the complex number. 1. |-4+3i|

OpenStudy (er.mohd.amir):

absolute value is =sqrt{(-4)^2+(3)^2}=sqrt(16+9)=sqrt(25)=ans

OpenStudy (joshoyen):

alright thanks!

OpenStudy (er.mohd.amir):

ans=sqrt(25)=5

OpenStudy (joshoyen):

Could you help me with one more?

OpenStudy (joshoyen):

How would you do |-10+8i|

OpenStudy (joshoyen):

I got sqaure root 164

OpenStudy (joshoyen):

what would the next step be

OpenStudy (joshoyen):

@Er.Mohd.AMIR

OpenStudy (anonymous):

Try this website: https://mathway.com/ It should help. : )

OpenStudy (joshoyen):

howd you get sqrt 2x9x9

OpenStudy (er.mohd.amir):

sorry 164=2*82

OpenStudy (joshoyen):

@Er.Mohd.AMIR oneeee more questoin haha sorry for |-10+6i| = 2 sqrt 17 ?

OpenStudy (er.mohd.amir):

sqrt 164 is ans or use calculator to find it.

OpenStudy (er.mohd.amir):

next quickly please

OpenStudy (joshoyen):

oh thats all but my paper says the answers is 2 sqrt 34

OpenStudy (joshoyen):

for |-10+6i|

OpenStudy (er.mohd.amir):

oh sqrt{(10)^2+(6)^2}=sqrt(100+36)=sqrt(136)=sqrt(2*2*34)=2 sqrt(34) is ans

OpenStudy (joshoyen):

couldn't you simplifity it more? to make it 2 sqrt 17 ?

OpenStudy (er.mohd.amir):

fine

OpenStudy (er.mohd.amir):

more

OpenStudy (joshoyen):

that was it, but couldn't you simplifity it evne more?

OpenStudy (joshoyen):

2 sqrt 34 to 2 sqrt 17

OpenStudy (er.mohd.amir):

no sqrt(x*x)=x if two number are same in sqrt one come out only

OpenStudy (joshoyen):

hmmm, i still don't get it

OpenStudy (joshoyen):

if which two numbers are the same?

OpenStudy (er.mohd.amir):

sqrt(136)=sqrt(2*2*34)=2 sqrt(34) s number two is same so out it 34=2*17 no number is same

OpenStudy (er.mohd.amir):

off

OpenStudy (joshoyen):

OHHH okay i see now, thank you!!

OpenStudy (denisaboichuk):

0